# Chain-of-Thought Prompting Obscures Hallucination Cues in Large Language Models: An Empirical Evaluation

### v1

@@ Abstract
Large Language Models (LLMs) often exhibit hallucinations, generating factually incorrect or semantically irrelevant content in response to prompts. Chain-of-Thought (CoT) prompting can mitigate hallucinations by encouraging step-by-step reasoning, but its impact on hallucination detection remains underexplored. To bridge this gap, we conduct a systematic empirical evaluation. We begin with a pilot experiment, revealing that CoT reasoning significantly affects the LLM’s internal states and token probability distributions. Building on this, we evaluate the impact of various CoT prompting methods on mainstream hallucination detection methods across both instruction-tuned and reasoning-oriented LLMs. Specifically, we examine three key dimensions: changes in hallucination score distributions, variations in detection accuracy, and shifts in detection confidence. Our findings show that while CoT prompting helps reduce hallucination frequency, it also tends to obscure critical signals used for detection, impairing the effectiveness of various detection methods. Our study highlights an overlooked trade-off in the use of reasoning. Code is publicly available at: https://anonymous.4open.science/r/cot-hallu-detect.
@@
Jiahao Cheng1  Tiancheng Su1  Jia Yuan1  Guoxiu He1 Jiawei Liu2Xinqi Tao3Jingwen Xie3Huaxia Li3 1 East China Normal University, Shanghai, China 2Wuhan University, Wuhan, China 3Xiaohsongshu Inc., Shanghai, China chengjiahao@stu.ecnu.edu.cn, gxhe@fem.ecnu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a range of natural language processing (NLP) tasks (Qin et al., 2024; Yan et al., 2024), showing a strong ability to follow instructions and answer questions (Zhao et al., 2024; Hadi et al., 2024). However, LLMs may also exhibit hallucinations. Specifically, they may produce factually incorrect or semantically irrelevant answers to a given prompt without indicating any lack of confidence (Ji et al., 2023; Yang et al., 2023; Hadi et al., 2024).
@@
Although some negative cases suggest that reasoning induced by CoT prompting may adversely affect hallucination detection, existing studies have yet to establish the generality of this phenomenon. To bridge this gap, we conduct a series of empirical evaluations. We begin with pilot experiments on three multiple-choice question answering (MCQA) tasks, including CommonSenseQA, ARC-Challenge, and MMLU, where we observe that CoT prompting significantly alters the probability distribution over the final answer tokens. Building on these findings, we perform a more extensive set of comparative experiments using two representative types of datasets: (1) complex fact-based question answering datasets, including TriviaQA, PopQA, HaluEval, and TruthfulQA; and (2) a summarization dataset requiring deep contextual understanding, named CNN/Daily Mail. We also incorporate four well-established, recently proposed hallucination detection methods to identify hallucinations. Furthermore, we evaluate three mainstream CoT prompting strategies across both instruction-tuned and reasoning-oriented LLMs, including LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, LLaMA-3.1-80B-Instruct, and DeepSeek-R1-Distill-Llama-8B. To comprehensively assess the impact of CoT prompting on hallucination detection, we examine three key dimensions: (1) shifts in the distribution of hallucination detection output scores; (2) changes in detection accuracy; and (3) variations in the confidence levels of the detection methods.
@@
Our contributions are threefold: (1) We conduct a series of experiments, including pilot and comprehensive evaluations, to empirically investigate the impact of CoT-induced reasoning on hallucination detection. (2) We reveal a double-edged effect of CoT prompting: while it enhances LLM performance, it simultaneously obscures common hallucination features. (3) Our results show that internal state-based and self-evaluation-based methods exhibit reduced robustness under CoT prompting.
@@ 2 Preliminary
CoT Prompting Methods
In this study, we focus on three representative methods: Zero-shot Chain of Thought (CoT, Kojima et al. (2024)), Least-to-Most (LtM, Zhou et al. (2023)), and Minimum Reasoning Path Prompting (MRPP, Chen et al. (2024c)). Detailed descriptions of these methods are provided in Appendix C.
Unless otherwise specified, all CoT prompting methods are implemented in a zero-shot configuration. This ensures that the LLM operates without access to task-specific examples, allowing us to isolate the effect of reasoning strategies. Following prior work, we adapt our objectives by tailoring prompt templates to suit each experimental setup.
Hallucination Detection Methods
We focus on hallucination detection methods that do not rely on external knowledge bases (e.g., fact-checking (Li et al., 2024)) or require additional fine-tuning and specialized training (e.g., ITI (Li et al., 2023b) and SAPLMA (Azaria and Mitchell, 2023)). This choice is motivated by the limited generalization ability of such methods across tasks and domains. Instead, we examine four main categories of hallucination detection methods: consistency-based methods (e.g., SelfCheckGPT-nli (Manakul et al., 2023)); internal-state-based methods (e.g., Perplexity, In-Context Sharpness (Chen et al., 2024d), and LLM-Check (Sriramanan et al., 2024)); self-evaluation-based methods (e.g., Verbalized Certainty (Kumar et al., 2024)); and hybrid methods (e.g., INSIDE (Chen et al., 2024a) and SelfCheckGPT-Prompt (Manakul et al., 2023)), which combine features from multiple categories. For more detailed descriptions of these methods, please refer to Appendix D.
@@ 3 Pilot Experiment and Observation
Recent research on multiple-choice question answering (MCQA) tasks (Kumar et al., 2024; Chen et al., 2024d; Quevedo et al., 2024) has revealed a strong correlation between the likelihood of a token being selected by an LLM and the correctness of the final answer. Building on these insights, we begin with a pilot experiment on token-level MCQA to examine how CoT-prompted reasoning influences the probability distribution over answer tokens. We pay particular attention to cases where incorrect answers or hallucinations persist despite the use of CoT prompting.
3.1 Pilot Experimental Setting
Datasets. We use three widely adopted MCQA datasets in our study. CommonsenseQA (validation split, 1,221 samples; Talmor et al. (2019)) is a benchmark designed to evaluate commonsense reasoning; we use its validation split since the test set lacks ground-truth answers. AI2 ARC (ARC-Challenge, test split, 1,172 samples; Boratko et al. (2018)) is a challenging dataset that focuses on science-related questions requiring advanced reasoning. MMLU (test split, 14,042 samples; Hendrycks et al. (2021)) is a large benchmark that spans a broad range of academic and professional domains, providing a comprehensive assessment of general knowledge and reasoning ability.
3.2 Observations
The experimental results are shown in Table 1. Overall, LLMs prompted with CoT consistently achieve higher accuracy across all datasets. For example, with Llama-3.1-8B-Instruct, CoT improves accuracy from 90.5, 90.16, and 80.79 to 97.67, 94.96, and 91.57 on the three datasets, respectively. However, the effectiveness of more advanced CoT variants is less consistent. While both LtM and MRPP employ multi-path reasoning, MRPP further incorporates confidence-based validation. As a result, LtM may suffer from ambiguity when faced with multiple plausible reasoning paths, achieving better performance than the base Llama but falling short of standard CoT. In contrast, MRPP outperforms even the basic CoT approach.
@@ 3.2 Observations
In addition, entropy consistently decreases across all LLMs after CoT prompting, suggesting that LLMs become more confident in their predictions, assigning higher probabilities to their selected answers. For instance, Llama’s entropy drops from 23.96, 26.33, and 46.94 to 6.63, 13.50, and 21.39, respectively, after applying CoT. Similar reductions are observed with other CoT strategies, reflecting a general improvement in model confidence.
The DeepSeek-R1-Distill-Llama-8B model, which is specifically optimized for reasoning abilities, performs substantially below expectations when reasoning is disabled. Its accuracy on the three datasets drops to 67.41, 59.26, and 54.62, respectively. When CoT is enabled, its performance improves significantly, though it still lags behind Llama in terms of accuracy. Interestingly, under CoT prompting, DeepSeek exhibits lower entropy than Llama across all datasets, suggesting greater confidence in its predictions, even though it produces more incorrect answers.
To better illustrate how confident LLMs are when producing hallucinated answers, i.e., incorrect predictions, we compare their token-level probability distributions before and after applying CoT prompting, for both all samples and hallucinated samples. The results are presented in Figure 2. Before CoT prompting, LLMs show relatively uniform probability distributions across samples; for hallucinated samples, the distributions are clearly left-skewed, indicating that incorrect answers tend to receive lower probabilities. In contrast, after CoT prompting, the distributions shift, and LLMs assign similarly high probabilities to both all and incorrect answers. This increased confidence, especially in wrong predictions, underscores the complex influence of CoT prompting on token-level probabilities, ultimately making hallucinations harder to detect.
@@
Previous studies have shown that the token probabilities produced by LLMs do not always reliably indicate hallucinations (Mahaut et al., 2024; Lin et al., 2022a; Chen et al., 2025). More robust signals can instead be extracted from the internal states of LLMs to better capture hallucination patterns. To explore this, we conduct a more comprehensive set of experiments using recent hallucination detection models, investigating whether CoT prompting influences the clarity of hallucination-related signals during the reasoning process.
4 Main Experiments
@@ 4.1 Pipeline
This section outlines the workflow used to assess whether CoT prompting interferes with hallucination detection, as illustrated in Figure 3. We begin by determining whether the LLM-generated response constitutes a hallucination based on the ground-truth response. Next, we compute a hallucination score using a detection model to predict whether the output is a hallucination. Finally, we examine the impact of CoT-generated reasoning on hallucination detection from three perspectives, including four evaluation metrics. Details are introduced below.
Correctness Measurement: For the responses generated by the LLM, we evaluate their alignment with the ground truth in the dataset based on text matching and semantic consistency. This evaluation determines whether a response constitutes a hallucination.
Hallucination Detection Score: During the response generation process, we apply four mainstream hallucination detection methods to compute the likelihood score of hallucination for each response. Based on these scores, we further predict whether a given response constitutes a hallucination.
Evaluation Dimensions:
Based on the hallucination labels and detection scores, we analyze the behavior of hallucination detection methods from surface-level patterns to deeper insights along three dimensions: score distribution, detection performance, and method confidence. Specifically: (1) changes in the distribution of hallucination scores are evaluated using a distribution consistency test; (2) detection performance is assessed via AUROC and the monotonic relationship; and (3) method confidence is measured using Expected Calibration Error (ECE). Further details are provided in Appendix E.
@@ 4.2 Experimental Setup
Datasets. We use four widely adopted QA datasets that require complex reasoning, along with one summarization dataset that demands strong in-context understanding. Specifically, we include TruthfulQA (Lin et al., 2022b), a dataset explicitly designed for hallucination detection, with a validation subset comprising 817 questions. In addition, TriviaQA (Joshi et al., 2017) is a large-scale reading comprehension dataset. We utilize its rc.wikipedia.nocontext and validation subsets. PopQA (Mallen et al., 2023) contains questions targeting a long-tail distribution of entities. We use its test subset in a closed-book setting. HaluEval (Li et al., 2023a) evaluates the factual accuracy of LLM-generated content by distinguishing between hallucinatory and accurate outputs. We use its QA subset of non-hallucinatory samples as a benchmark for correctness. For summarization, we employ the CNN/Daily Mail dataset Hermann et al. (2015), where the LLM is required to generate a summary based on a given news article. This task emphasizes the need for faithful summarization without deviation from the source content.
LLMs. Since some hallucination detection methods require access to the internal states of LLMs, we conduct our experiments using open-source LLMs. To ensure diversity across model families, parameter scales, and training objectives, we select four LLMs, including Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Llama-3.1-80B-Instruct, and DeepSeek-R1-Distill-Llama-8B.
The implementation details of the experiment can be found in Appendix F.
@@ 5.1 Performance Comparison of CoT Methods Across Different Datasets
As shown in Table 2, consistent with the results of our pilot experiments, the CoT prompting methods significantly enhance the performance of LLMs across various tasks. By encouraging LLMs to explicitly generate intermediate reasoning steps, CoT prompting helps activate their latent knowledge and reasoning capabilities, leading to more accurate responses.
Notably, LLM performance rankings vary between the two evaluation approaches, textual matching and semantic consistency, highlighting the challenge posed by the diverse linguistic expressions in text generation. While evaluation methodology is not the primary focus of this study, to mitigate ambiguity, we adopt both metrics in subsequent experiments to assess response correctness, which serves as the ground truth for hallucination detection.

### v3

@@ Abstract
Large Language Models (LLMs) often exhibit hallucinations, generating factually incorrect or semantically irrelevant content in response to prompts. Chain-of-Thought (CoT) prompting can mitigate hallucinations by encouraging step-by-step reasoning, but its impact on hallucination detection remains underexplored. To bridge this gap, we conduct a systematic empirical evaluation. We begin with a pilot experiment, revealing that CoT reasoning significantly affects the LLM’s internal states and token probability distributions. Building on this, we evaluate the impact of various CoT prompting methods on mainstream hallucination detection methods across both instruction-tuned and reasoning-oriented LLMs. Specifically, we examine three key dimensions: changes in hallucination score distributions, variations in detection accuracy, and shifts in detection confidence. Our findings show that while CoT prompting helps reduce hallucination frequency, it also tends to obscure critical signals used for detection, impairing the effectiveness of various detection methods. Our study highlights an overlooked trade-off in the use of reasoning. Code is publicly available at: https://github.com/ECNU-Text-Computing/cot-hallu-detect.
@@
Jiahao Cheng1  Tiancheng Su1  Jia Yuan1  Guoxiu He1thanks: Corresponding author. Jiawei Liu2Xinqi Tao3Jingwen Xie3Huaxia Li3 1 East China Normal University, Shanghai, China 2Wuhan University, Wuhan, China 3Xiaohongshu Inc., Shanghai, China chengjiahao@stu.ecnu.edu.cn, gxhe@fem.ecnu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a range of natural language processing (NLP) tasks (Qin et al., 2024; Yan et al., 2024; Song et al., 2025), showing a strong ability to follow instructions and answer questions (Zhao et al., 2024; Hadi et al., 2024). However, LLMs may also exhibit hallucinations. Specifically, they may produce factually incorrect or semantically irrelevant answers to a given prompt without indicating any lack of confidence (Ji et al., 2023; Yang et al., 2023; Hadi et al., 2024).
@@
Although several counterexamples indicate that reasoning induced by CoT prompting may negatively impact hallucination detection, the generality of this effect has not been established in prior research. To bridge this gap, we conduct a series of empirical evaluations. We start with pilot experiments on three multiple-choice question answering (MCQA) tasks: CommonSenseQA, ARC-Challenge, and MMLU. In these experiments, we observe that CoT prompting significantly alters the probability distribution of the final answer tokens. Building on these findings, we perform a more extensive set of comparative experiments using two representative types of datasets: (1) complex fact-based question answering datasets, including TriviaQA, PopQA, HaluEval, and TruthfulQA; and (2) a summarization dataset requiring deep contextual understanding, named CNN/Daily Mail. We also incorporate four well-established, recently proposed hallucination detection methods to identify hallucinations. Furthermore, we evaluate three mainstream CoT prompting strategies across both instruction-tuned and reasoning-oriented LLMs, including LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, LLaMA-3.1-70B-Instruct, and DeepSeek-R1-Distill-Llama-8B. To comprehensively assess the impact of CoT prompting on hallucination detection, we examine three key dimensions: (1) shifts in the distribution of hallucination detection output scores; (2) changes in detection accuracy; and (3) variations in the confidence levels of the detection methods.
@@
Our contributions are threefold: (1) We conduct a series of experiments, including pilot and comprehensive evaluations, to empirically investigate the impact of CoT-induced reasoning on hallucination detection. (2) We reveal a double-edged effect of CoT prompting: while it enhances LLM performance, it simultaneously obscures common hallucination features. (3) Our results show that internal-state-based and self-evaluation-based methods exhibit reduced robustness under CoT prompting.
@@ 2 Preliminary
CoT Prompting Methods.
In this study, we focus on three representative methods: Zero-shot Chain of Thought (CoT, Kojima et al. (2024)), Least-to-Most (LtM, Zhou et al. (2023)), and Minimum Reasoning Path Prompting (MRPP, Chen et al. (2024c)). Further details and formal definitions of these methods are provided in Appendix C.
Unless otherwise specified, all CoT prompting methods are implemented in a zero-shot configuration. This ensures that the LLM operates without access to task-specific examples, allowing us to isolate the effects of different reasoning strategies. Following prior work, we design prompt templates tailored to each experimental configuration to align with specific task requirements.
Hallucination Detection Methods.
We focus on hallucination detection methods that do not rely on external knowledge bases (e.g., fact-checking (Li et al., 2024)) or require additional fine-tuning and specialized training (e.g., ITI (Li et al., 2023b) and SAPLMA (Azaria and Mitchell, 2023)). This choice is motivated by the limited generalization ability of such methods across tasks and domains. Instead, we examine four main categories of hallucination detection methods: consistency-based methods (e.g., SelfCheckGPT-nli (Manakul et al., 2023)); internal-state-based methods (e.g., Perplexity, In-Context Sharpness (Chen et al., 2024d), and LLM-Check (Sriramanan et al., 2024)); self-evaluation-based methods (e.g., Verbalized Certainty (Kumar et al., 2024)); and hybrid methods (e.g., INSIDE (Chen et al., 2024a) and SelfCheckGPT-Prompt (Manakul et al., 2023)), which combine features from multiple categories. For comprehensive descriptions and further details of these methods, see Appendix D.
@@ 3 Pilot Experiment and Observation
Recent studies on multiple-choice question answering (MCQA) tasks (Kumar et al., 2024; Chen et al., 2024d; Quevedo et al., 2024) have identified a strong correlation between the probability assigned by LLMs to answer tokens and the correctness of their final predictions. In these datasets, which are primarily composed of factual questions, we adopt the common practice of treating answer incorrectness as a proxy for hallucination labels (Kumar et al., 2024). Motivated by these insights, we perform a pilot experiment to analyze, at the token level, how reasoning with chain-of-thought (CoT) prompting shapes the distribution of probabilities over possible answer tokens. Our analysis pays special attention to instances where CoT prompting fails to eliminate incorrect (i.e., hallucinated) answers.
3.1 Pilot Experimental Setting
Datasets. We use three widely adopted MCQA datasets in our study. CommonsenseQA (validation split, 1,221 samples; Talmor et al. (2019)) is a benchmark designed to evaluate commonsense reasoning; we use the validation split since ground-truth answers are unavailable for its test set. AI2 ARC (ARC-Challenge, test split, 1,172 samples; Boratko et al. (2018)) is a challenging dataset that focuses on science-related questions requiring advanced reasoning. MMLU (test split, 14,042 samples; Hendrycks et al. (2021)) is a large benchmark that spans a broad range of academic and professional domains, providing a comprehensive assessment of general knowledge and reasoning ability.
@@ 3.1 Pilot Experimental Setting
3.2 Observations
The experimental results are shown in Table 1. Overall, LLMs prompted with CoT consistently achieve higher accuracy across all datasets. For example, with Llama-3.1-8B-Instruct, CoT improves accuracy (%) from 90.5, 90.16, and 80.79 to 97.67, 94.96, and 91.57 on the three datasets, respectively. However, the effectiveness of more advanced CoT variants is less consistent. While both LtM and MRPP employ multi-path reasoning, MRPP further incorporates confidence-based validation. As a result, LtM may suffer from ambiguity when faced with multiple plausible reasoning paths, achieving better performance than the base Llama but falling short of standard CoT. In contrast, MRPP outperforms even the basic CoT approach.
In addition, after applying CoT prompting, entropy consistently decreases across all LLMs. This suggests that LLMs become more confident in their predictions by assigning higher probabilities to their selected answers. For instance, Llama’s entropy drops from 23.96, 26.33, and 46.94 to 6.63, 13.50, and 21.39, respectively, after applying CoT. Similar reductions are observed with other CoT strategies, reflecting a general improvement in model confidence.
The DeepSeek-R1-Distill-Llama-8B model, which is specifically optimized for reasoning abilities, performs substantially below expectations when reasoning is disabled. Its accuracy on the three datasets drops to 67.41, 59.26, and 54.62, respectively. When CoT is enabled, its performance improves significantly, though it still lags behind Llama in terms of accuracy. Interestingly, under CoT prompting, DeepSeek exhibits lower entropy than Llama across all datasets. This suggests that DeepSeek is more confident in its predictions, even though it produces more incorrect answers.
To better illustrate how confident LLMs are when producing hallucinated answers (i.e., incorrect predictions), we compare their token-level probability distributions before and after applying CoT prompting. This comparison is conducted for both all samples and hallucinated samples. The results are presented in Figure 2. Before CoT prompting, LLMs show relatively uniform probability distributions across samples. For hallucinated samples, the distributions are clearly left-skewed, indicating that incorrect answers tend to receive lower probabilities. In contrast, after CoT prompting, the probability distributions shift such that LLMs assign similarly high probabilities to both correct and incorrect answers. This increased confidence, especially in wrong predictions, underscores the complex influence of CoT prompting on token-level probabilities, ultimately making hallucinations harder to detect.
Previous studies have shown that token probabilities produced by LLMs do not always reliably indicate the presence of hallucinations (Mahaut et al., 2024; Lin et al., 2022a; Chen et al., 2025). Instead, more robust signals can be extracted from the internal states of LLMs to better capture hallucination patterns. To explore this, we conduct a more comprehensive set of experiments using recent hallucination detection models. We investigate whether CoT prompting influences the clarity of hallucination-related signals during the reasoning process.
4 Main Experiments
@@ 4.1 Pipeline
This section outlines the workflow used to assess whether CoT prompting interferes with hallucination detection, as illustrated in Figure 3. We begin by determining whether the LLM-generated response constitutes a hallucination based on the ground-truth response. Next, we use a detection model to compute a hallucination score for each output, which is then used to predict whether the output is a hallucination. Finally, we examine the impact of CoT-generated reasoning on hallucination detection from three perspectives. These include four evaluation metrics.
Correctness Measurement. For the responses generated by the LLM, we evaluate their alignment with the ground truth in the dataset based on text matching and semantic consistency. This evaluation determines whether a response constitutes a hallucination.
Hallucination Detection Score. During the response generation process, we apply four mainstream hallucination detection methods to compute the likelihood score of hallucination for each response. Based on these scores, we further predict whether a given response constitutes a hallucination.
Evaluation Dimensions. Based on the hallucination labels and detection scores, we analyze the behavior of hallucination detection methods along three dimensions: score distribution, detection performance, and method confidence. These dimensions range from surface-level statistical patterns to deeper model-level insights. Specifically: (1) changes in the distribution of hallucination scores are evaluated using a distribution consistency test; (2) detection performance is assessed via AUROC and the monotonic relationship; and (3) method confidence is measured using Expected Calibration Error (ECE). Further details are provided in Appendix E.
@@ 4.2 Experimental Setup
Datasets. We use four widely adopted QA datasets that require complex reasoning and one summarization dataset. The summarization task demands strong in-context understanding. Specifically, we include TruthfulQA (Lin et al., 2022b), a dataset explicitly designed for hallucination detection. We use its validation subset, which comprises 817 questions. In addition, TriviaQA (Joshi et al., 2017) is a large-scale reading comprehension dataset. We utilize its rc.wikipedia.nocontext split and validation subset to support our experiments. PopQA (Mallen et al., 2023) contains questions targeting a long-tail distribution of entities. We use its test subset in a closed-book setting to avoid external knowledge access during evaluation. HaluEval (Li et al., 2023a) evaluates the factual accuracy of LLM-generated content by distinguishing between hallucinatory and accurate outputs. We use its QA subset of non-hallucinatory samples as a benchmark for correctness. For summarization, we employ the CNN/Daily Mail dataset Hermann et al. (2015), where the LLM is required to generate a summary based on a given news article. This task emphasizes faithful summarization, requiring that generated summaries do not deviate from the source content.
LLMs. Some hallucination detection methods require access to the internal states of LLMs. Thus, we conduct our experiments using open-source LLMs. To ensure diversity across model families, parameter scales, and training objectives, we select four LLMs, including Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Llama-3.1-80B-Instruct, and DeepSeek-R1-Distill-Llama-8B.
Implementation details of the experiments are provided in Appendix F.
@@ 5 Results and analysis
This section presents a comprehensive analysis of the experimental results, focusing on three key aspects: (1) the performance enhancement of LLMs through CoT prompting across diverse tasks and datasets; (2) the impact of CoT prompting on the internal mechanisms of LLMs, as revealed by the distribution shifts in hallucination detection scores; and (3) the consequential effects on the performance and reliability of hallucination detection methods. Our findings demonstrate that while CoT prompting significantly improves output quality, it also alters the models’ internal representations and confidence calibration. As a result, this ultimately complicates the hallucination detection process.
We validate our automated metrics by having two annotators manually annotate 150 samples, achieving a Cohen’s Kappa of 0.863. The Spearman’s $\rho$ between our automated scores and human annotations is 0.690 for Exact Match, 0.656 for ROUGE-L F1, and 0.683 for semantic similarity. These results confirm that our automated evaluations reliably reflect human judgment.
@@ 5.1 Performance Comparison of CoT Methods Across Different Datasets
As shown in Table 2, the CoT prompting methods significantly enhance the performance of LLMs across various tasks. This is consistent with the results of our pilot experiments. By encouraging LLMs to explicitly generate intermediate reasoning steps, CoT prompting helps activate their latent knowledge and reasoning capabilities, leading to more accurate responses.
Notably, LLM performance rankings vary between the two evaluation approaches, namely textual matching and semantic consistency. This highlights the challenge posed by diverse linguistic expressions in text generation. While evaluation methodology is not the primary focus of this study, we adopt both metrics in subsequent experiments to assess response correctness. This approach helps to mitigate ambiguity, and the resulting correctness serves as the ground truth for hallucination detection.